Ace Industrial Machines issued 190,000 zero coupon bonds 7 years ago. The bonds originally had 30 years to maturity with a yield to maturity of 6.6 percent. Interest rates have recently decreased, and the bonds now have a yield to maturity of 5.7 percent. The bonds have a par value of $2,000. If the company has a $88.2 million market value of equity, what weight should it use for debt when calculating the cost of capital? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1

Price of Bond= 2,000/(1.057)^23

Price of Bond = 558.86

Value of debt = 558.86(190,000) = $106.18 million

Weight of Debt = 106.18/106.18 + 88.2)

Weight of Debt = 54.62%